OpenStudy (anonymous):

Question:

OpenStudy (anonymous):

\[a+4\sqrt{3a ^{3}+10a ^{2}-5a+12}\]

OpenStudy (vishweshshrimali5):

Ok what do we have to do ?

OpenStudy (anonymous):

That is what I am wondering. :) The problem is above

OpenStudy (vishweshshrimali5):

Okay see: \[\large{3a^3 + 10a^2 - 5a + 12 = 3a^2 (a+4) - 2a(a+4) + 3(a+4)}\] \[\large{=(a+4)(3a^2 - 2a + 3)}\]

OpenStudy (marissalovescats):

I think she has to do and show long division

OpenStudy (anonymous):

yes I do :)

OpenStudy (anonymous):

I am a little unsure how to start it still

OpenStudy (vishweshshrimali5):

 Continue after this

OpenStudy (anonymous):